# Entropy Stability and Spectral Concentration under Convex Block Constraints

### v1

@@ 7.1 Pinsker and Csiszár–Kullback inequalities
By contrast, Theorem 4.1 concerns the entropy functional itself as a variational object over a constraint set. No reference state is fixed in advance; rather, the entropy minimizer emerges intrinsically from the geometry of the constraint set.
Remark 7.1.
Pinsker’s inequality provides lower bounds on relative entropy, whereas Theorem 4.1 provides lower bounds on entropy gaps within constrained families.
7.2 Strong convexity and entropy
@@
It is well known that the von Neumann entropy is concave but not strongly concave on the full state space. In particular, there is no global constant $c>0$ such that
$S\!\left(\tfrac{1}{2}(\rho+\sigma)\right)\;\geq\;\tfrac{1}{2}S(\rho)+\tfrac{1}{2}S(\sigma)+c\|\rho-\sigma\|_{1}^{2}$
for all states $\rho,\sigma$.
The novelty of Theorem 4.1 lies in showing that strong convexity emerges after restriction to block-convex constraint sets. The constant $c$ depends only on the block geometry, not on the ambient dimension.
Proposition 7.2 (Conditional strong convexity).
Let $\mathcal{C}$ be block-convex. Then S-S is strongly convex on $\mathcal{C}$:
$S(\rho)-S(\sigma)\;\geq\;c\,\|\rho-\sigma\|_{1}^{2},\qquad\rho\in\mathcal{C},$
where $\sigma$ is the entropy minimizer.
This phenomenon does not hold on general convex subsets of $\mathcal{S}(\mathcal{H})$.
7.3 Log-Sobolev inequalities
Log-Sobolev inequalities play a central role in entropy decay and mixing theory. A typical quantum log-Sobolev inequality has the form
$S(\rho\|\sigma)\;\leq\;C\,\mathcal{E}(\rho),$
where $\mathcal{E}$ is a Dirichlet form associated with a quantum Markov semigroup.
Our stability inequality is orthogonal in nature:
  • No dynamics or semigroup is involved
  • No reference equilibrium state is fixed
  • The inequality is purely static and geometric
Nevertheless, one may interpret Theorem 4.1 as a static precursor to log-Sobolev behavior restricted to symmetry sectors.
7.4 Majorization and entropy minimizers
Entropy minimization under linear constraints is closely related to majorization theory. Given a fixed block structure, entropy is minimized by states whose eigenvalue distribution is as uneven as permitted.
Proposition 7.3 (Majorization minimality).
Let $\mathcal{C}$ be block-convex and let $\sigma$ minimize entropy on $\mathcal{C}$. Then for any $\rho\in\mathcal{C}$, the spectrum of $\rho$ majorizes the spectrum of $\sigma$.
Proof.
Entropy is Schur-concave. The minimizer must therefore lie on the minimal face of the majorization order within $\mathcal{C}$. ∎
The stability theorem quantifies how strictly this majorization ordering controls distance in trace norm.
7.5 Why the present inequality is new
We emphasize the following distinctions:
  1. 1.
    The inequality is intrinsic: no reference state is fixed.
  2. 2.
    The constant depends only on block geometry, not on dimension.
  3. 3.
    The inequality is second-order stable, not merely qualitative.
  4. 4.
    The setting is infinite-dimensional but finite-block.
We are not aware of prior results establishing quantitative entropy stability for entropy minimizers under general block-convex constraints of the type considered here.
7.6 Minimal hypotheses
The proof of Theorem 4.1 requires only:
  • Block-diagonal structure
  • Convexity of the constraint set
  • Lower semicontinuity of entropy
No spectral gap, no finite-dimensionality assumption, and no arithmetic structure are required.
7.7 Positioning within the literature
The present work complements classical results of:
  • Pinsker, Csiszár, Kullback (relative entropy)
  • Lieb–Ruskai (strong subadditivity)
  • Carlen–Maas (entropy and geometry)
by addressing a distinct but natural question: how rigid is entropy minimization under structured constraints?
@@ 8 Conclusion and Outlook
8.1 Summary of results
This paper establishes a general stability principle for entropy minimizers under structured convex constraints. The main result, Theorem 4.1, shows that when a convex constraint set admits a finite block decomposition, entropy minimization is quantitatively rigid: deviations from the minimizer incur a quadratic entropy penalty controlled by trace distance.
The stability constant is independent of the ambient dimension, provided the block structure and constraint geometry are fixed.”, on arithmetic structure, or on dynamical assumptions. The stability constant depends only on the geometry of the block decomposition and not on the underlying system size.
Several refinements are established:
  • sharpness of the inequality and characterization of extremal states,
  • equivalent formulations in terms of purity and effective dimension,
  • majorization constraints governing admissible spectral distributions,
  • stability under affine and symmetry-preserving perturbations.
These results identify entropy as a strictly stable variational functional once symmetry-induced degeneracies are properly factored out.
8.2 Conceptual significance
From a conceptual perspective, the main contribution is the identification of a mechanism by which strong convexity of entropy emerges conditionally, despite its absence at the global level. The block structure plays a role analogous to symmetry reduction in classical mechanics: it removes flat directions and reveals intrinsic curvature of the entropy functional.
This viewpoint aligns entropy minimization with classical stability theory in convex optimization and provides a rigorous foundation for treating entropy as an energy-like quantity within constrained families.
8.3 Relation to previous work
While entropy inequalities have been extensively studied, most classical results focus either on relative entropy between two fixed states or on dynamical entropy decay. The present work addresses a distinct question: quantitative rigidity of entropy minimizers under static convex constraints.
To our knowledge, no prior result establishes a universal stability inequality of the form proved here for general block-convex sets. The analysis complements existing literature on entropy, majorization, and convex geometry without overlapping with known results.
8.4 Limitations
All results in this paper are proved under the standing assumption that the Hilbert space $\mathcal{H}$ is finite-dimensional. This assumption is essential for compactness and for uniform control of entropy and trace-norm estimates.
It is plausible that analogous stability phenomena persist in certain infinite-dimensional settings with a finite number of blocks and appropriate tightness or compactness conditions. However, extending the present arguments to such settings would require substantial additional work and is beyond the scope of this paper.
8.5 Future directions
The results suggest several natural directions for further investigation:
  1. 1.
    Optimal constants. Determining sharp stability constants in terms of geometric invariants of block decompositions.
  2. 2.
    Infinite block limits. Extending stability to settings with controlled infinite block families.
  3. 3.
    Dynamical extensions. Connecting static entropy stability to entropy production along quantum Markov semigroups.
  4. 4.
    Non-commutative constraints. Investigating stability under constraints generated by non-abelian algebras.
  5. 5.
    Applications. Exploring implications in quantum information, statistical mechanics, and spectral theory.
These directions suggest that entropy stability under symmetry and convexity constraints may constitute a broader principle with applications beyond the specific setting considered here.
8.6 Final remarks
The present work isolates a simple but robust mechanism underlying entropy rigidity. By stripping away arithmetic and dynamical language, the results are placed firmly within functional analysis and convex geometry.
We hope that this perspective will prove useful both as a technical tool and as a conceptual guide for future studies of entropy-driven phenomena.
Appendix A Appendix: Standard Entropy Facts
@@ A.1 Von Neumann entropy
Lemma A.1 (Lower semicontinuity).
@@
Proof.
This follows from Fatou’s lemma applied to the spectral decomposition of $\rho$. See [Wehrl1978, Simon1979]. ∎
A.2 Entropy of block-diagonal states
@@
Lemma A.2 (Entropy decomposition).
@@
Proof.
Since the blocks act on orthogonal subspaces, the spectrum of $\rho$ is the union of the spectra of the operators $\mu(i)\rho_{i}$. The claim follows by direct calculation. ∎
A.3 Elementary entropy bounds
@@
Lemma A.3 (Entropy upper bound).
@@
Proof.
This is classical. The Shannon entropy is strictly concave on the simplex $\Delta_{r}$ and is maximized uniquely at the uniform distribution. ∎
Lemma A.4 (Maximum mass lower bound).
@@
which implies $M\geq e^{-H(\mu)}$. ∎
Remark A.5.
The bound in Lemma A.4 is sharp, with equality attained for distributions supported on a single point.
Lemma A.6 (Support-size refined bound).
@@
Proof.
By Lemma A.4, $\max_{i}\mu_{i}\geq e^{-H(\mu)}$. Since $\sum_{i}\mu_{i}=1$ and there are at most $r$ nonzero terms,
@@
Remark A.7.
Lemma A.6 is weaker than Lemma A.4 for small entropy, but is convenient when an explicit dependence on support size is needed.
A.4 Trace norm and block structure
Lemma A.8 (Blockwise trace norm).
@@ A.4 Trace norm and block structure
Proof.
The operators act on orthogonal subspaces, and the trace norm is additive over orthogonal direct sums. ∎
A.5 Pinsker inequality
Lemma A.9 (Pinsker).
@@ A.5 Pinsker inequality
Proof.
See [Csiszar1967, Petz2001]. ∎
A.6 Use within the paper
The results above are used as follows:
  • Lemma A.2 underlies the entropy decomposition in Sections 4 and 5.
  • Lemma A.6 is used in bounding entropy gaps.
@@ A.6 Use within the paper
  • Lemma A.8 is used to control trace distances between block-diagonal states.
  • Lemma A.9 is used for relative-entropy comparisons.
No other entropy inequalities are invoked implicitly.

### v2

@@ 7.1 Relation to Pinsker, continuity, and functional inequalities
These inequalities are complementary in spirit to the present work: they control either relative entropy with respect to a fixed reference state or uniform continuity of entropy on the full state space. By contrast, Theorem 4.1 treats the entropy functional itself as a variational object over a structured constraint set, with no reference state fixed a priori; the entropy minimizer instead emerges intrinsically from the geometry of the constraint.
Log-Sobolev inequalities relate entropy to Dirichlet forms associated with Markov semigroups and play a central role in entropy decay and mixing theory [KastoryanoTemme2013, CarlenMaas2017, Brannan2022]. The present stability estimate is orthogonal in nature: it is purely static and geometric, involving neither dynamics nor a prescribed equilibrium state. Related perspectives on entropy inequalities and stability appear in classical convex analysis and information theory; see, e.g., [Kullback1959, Rockafellar1970, BonnansShapiro2000].
7.2 Quadratic growth and nonstandard features
@@
It is well known that the von Neumann entropy is concave but not strongly concave on the full state space. Consequently, no global quadratic growth inequality can hold without further structural assumptions.
The nonstandard feature of Theorem 4.1 is that quadratic entropy–distance stability emerges locally near constrained minimizers after restriction to block-convex constraint sets. The entropy functional lacks global strong convexity, and the admissible set is stratified; quadratic growth therefore does not hold a priori, but arises only after confinement to a fixed-support stratum determined by the constraint geometry.
From an optimization viewpoint, the theorem establishes a quadratic growth (error bound) property for entropy minimization on a structured convex set, with constants depending only on the geometry of the constraint and not on the ambient Hilbert space dimension.
This setting differs fundamentally from classical majorization or continuity results, which describe qualitative extremality but do not provide quantitative stability estimates near entropy minimizers.
8 Computational remarks
The stability constant $C$ in Theorem 4.1 can be made explicit for concrete constraint sets by combining marginal and conditional second-order information.
Marginal contribution.
The marginal constant $c_{1}$ from Lemma 4.3 is determined by the strict concavity of Shannon entropy on faces of the marginal polytope $\Pi$. For an extreme point $q\in\operatorname{ext}(\Pi)$, let $F$ denote the minimal face containing $q$. The Hessian of $H(p)=-\sum_{i}p_{i}\log p_{i}$ restricted to the tangent space of $F$ at $q$ is negative definite, with eigenvalues given by $1/q_{i}$ along admissible directions. The constant $c_{1}$ is obtained by minimizing the smallest absolute eigenvalue over all such faces $F$. For polytopes specified by a finite block structure, this quantity can be computed explicitly or bounded from below.
Conditional contribution.
The conditional constant $c_{2}$ from Lemma 3.8 depends on the Hessian of von Neumann entropy at the conditional minimizers $\sigma_{i}\in\mathcal{C}_{i}$. Under the confining hypothesis, the entropy Hessian is uniformly positive on admissible tangent directions, with a lower bound $m_{i}>0$ determined by the local spectral gap and geometry of $\mathcal{C}_{i}$. Setting $c_{2}=\min_{i:q_{i}>0}m_{i}$, the overall stability constant takes the form
$C=\min\!\left\{c_{1},\;\frac{c_{2}}{2d_{\max}}\right\},$
where $d_{\max}$ accounts for norm equivalence on conditional subspaces.
Remarks on computation.
For constraint sets of moderate size, the constants $c_{1}$ and $c_{2}$ can be evaluated numerically using standard eigenvalue routines. In higher-dimensional settings, computable lower bounds follow from symmetry, block structure, or sampling of admissible tangent directions. These considerations confirm that the stability constant is not merely existential, but can be estimated effectively from the geometry of the constraint. Related computational aspects of entropy and convex optimization are discussed in [Watrous2018, Wilde2017].
@@ 9 Conclusion and outlook
We have established a second-order stability principle for entropy minimization under block-separable convex constraints. The main result, Theorem 4.1, shows that under a confining (fixed-support) hypothesis, the entropy gap controls the squared trace-norm distance to the set of minimizers, with explicit constants determined by the geometry of the constraint.
The analysis reveals a clear geometric mechanism underlying quadratic stability. Although the von Neumann entropy lacks global strong convexity, genuine second-order rigidity emerges locally near constrained minimizers once admissible perturbations are confined to fixed-support strata. The block-separable structure allows the stability problem to decompose into independent marginal and conditional components, whose contributions are governed by the Hessian curvature of Shannon and von Neumann entropy, respectively.
We further show that the quadratic exponent is optimal: even in classical (commutative) settings, no uniform linear or subquadratic stability bound can hold near entropy minimizers. This optimality confirms that the stability mechanism identified here is sharp and intrinsic to the geometry of the constraint.
All results are obtained in finite dimension, where compactness and norm equivalence ensure uniform control of stability constants. Extensions to infinite-dimensional settings or to more general (non block-separable) constraints would require substantially different techniques and are left for future work.
Overall, this work provides a complete and sharp second-order stability theory for entropy minimization under structured convex constraints, with both conceptual and practical implications for optimization and information theory.
Appendix A Domain of validity of quadratic entropy–distance stability
This appendix describes geometric regimes that are explicitly excluded from Theorem 4.1 and explains why a purely quadratic entropy–distance stability estimate cannot hold outside the confining (fixed-support) hypothesis.
Its purpose is not to extend the main result, but to delineate sharply the boundary of its applicability and to show that the failure of quadratic stability in boundary–transversal regimes is structural and unavoidable. This clarification renders the scope of the entropy-based arguments used throughout the paper logically complete and internally consistent.
A.1 Quadratic stability under fixed-support (confining) constraints
Let $\mathcal{C}\subset\mathcal{D}(\mathcal{H})$ be a convex constraint set and let $\sigma\in\mathcal{C}$ be an entropy minimizer. Denote by $P=\operatorname{supp}(\sigma)$ the support projection of $\sigma$.
Lemma A.1 (Quadratic entropy stability in the confining regime).
Assume that there exists a neighbourhood $U$ of $\sigma$ in $\mathcal{C}$ such that
$\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)\qquad\text{for all }\rho\in U.$
Equivalently, admissible perturbations are confined to the fixed-support stratum $\mathcal{D}(P\mathcal{H})$. Then there exist constants $c,C>0$ such that, for all $\rho\in U$,
$c\,\|\rho-\sigma\|_{1}^{2}\;\leq\;S(\rho)-S(\sigma)\;\leq\;C\,\|\rho-\sigma\|_{1}^{2}.$
Proof A.2 (Proof sketch).
On the fixed-support stratum $\mathcal{D}(P\mathcal{H})$, the von Neumann entropy $S$ is a $C^{2}$ functional. Since $\sigma$ is full rank on its support, that is, $P\sigma P\succ 0$ as an operator on $P\mathcal{H}$, the Hessian of f:=Sf:=-S at $\sigma$ is strictly positive on the admissible tangent directions compatible with the constraint. A second-order Taylor expansion therefore yields the stated local quadratic upper and lower bounds in trace norm.
A.2 Failure outside the confining regime
The conclusion of Lemma A.1 relies crucially on the fixed-support (confining) hypothesis. If this assumption fails-specifically, if the entropy minimizer $\sigma$ is rank-deficient and the constraint allows admissible perturbations to explore directions transversal to the support-then the entropy Hessian degenerates at $\sigma$. In such boundary-touching, transversal regimes, the quadratic estimate above is no longer valid.
In that case, the optimal entropy–distance modulus acquires an unavoidable logarithmic correction, and one enters the quadratic–logarithmic rigidity regime analysed in the main body of this work. This distinction explains why purely quadratic entropy-based stability estimates cannot hold uniformly across all symmetry-restricted settings.
A.3 Relation to the main results
Lemma A.1 isolates the geometric regime in which classical quadratic entropy–distance estimates are valid. The main rigidity theorem addresses the complementary geometric regime in which the minimizer remains confined, while boundary–transversal regimes are treated elsewhere.
Together, these results provide a complete geometric picture of entropy stability near constrained minimizers.
Appendix B Standard Entropy Facts
@@ B.1 Von Neumann entropy
Lemma B.1 (Lower semicontinuity).
@@
Proof B.2.
This follows from Fatou’s lemma applied to the spectral decomposition of $\rho$. See [Wehrl1978, Simon1979].
B.2 Entropy of block-diagonal states
@@
Lemma B.3 (Entropy decomposition).
@@
Proof B.4.
Since the blocks act on orthogonal subspaces, the spectrum of $\rho$ is the union of the spectra of the operators $\mu(i)\rho_{i}$. The claim follows by direct calculation.
B.3 Elementary entropy bounds
@@
Lemma B.5 (Entropy upper bound).
@@
Proof B.6.
This is classical. The Shannon entropy is strictly concave on the simplex $\Delta_{r}$ and is maximized uniquely at the uniform distribution.
Lemma B.7 (Maximum mass lower bound).
@@
which implies $M\geq e^{-H(\mu)}$.
Remark B.9.
The bound in Lemma B.7 is sharp, with equality attained for distributions supported on a single point.
Lemma B.10 (Support-size refined bound).
@@
The purely quadratic estimate used here is valid only in the fixed-support (confining) regime; see Appendix A for a precise geometric formulation and proof.
Proof B.11.
By Lemma B.7, $\max_{i}\mu_{i}\geq e^{-H(\mu)}$. Since $\sum_{i}\mu_{i}=1$ and there are at most $r$ nonzero terms,
@@
Remark B.12.
Lemma B.10 is weaker than Lemma B.7 for small entropy, but is convenient when an explicit dependence on support size is needed.
B.4 Trace norm and block structure
Lemma B.13 (Blockwise trace norm).
@@ B.4 Trace norm and block structure
Proof B.14.
The operators act on orthogonal subspaces, and the trace norm is additive over orthogonal direct sums.
B.5 Pinsker inequality
Lemma B.15 (Pinsker).
@@ B.5 Pinsker inequality
Proof B.16.
See [Csiszar1967, Petz2001].
B.6 Use within this work
These results are used explicitly as quantitative entropy inequalities in the proofs of the main theorems:
  • Lemma B.3 underlies the entropy decomposition in Sections 4 and 5.
  • Lemma B.10 is used to bound entropy gaps.
@@ B.6 Use within this work
  • Lemma B.13 is used to control trace distances between block-diagonal states.
  • Lemma B.15 is used for relative-entropy comparisons.
No further nontrivial entropy inequalities beyond standard properties are invoked at the level of quantitative estimates. The boundary–transversal regime, which leads to logarithmic corrections, is treated separately and will be discussed elsewhere.
@@
Conflict of Interest
The author declares that there is no conflict of interest regarding the publication of this article.
Funding
The author received no external funding.
Code availability
Not applicable.